# Training-Free Message Passing for Learning on Hypergraphs

### v4

@@ 7 Conclusion
In this paper, we propose a novel model called TF-HNN. The key innovation of TF-HNN is an original training-free message passing module (TF-MP-Module) tailored for data on hypergraphs. We present both theoretical and empirical evidence demonstrating that TF-HNN can efficiently and effectively address hypergraph-related downstream tasks. To our knowledge, TF-HNN is the first model to shift the integration of hypergraph structural information from the model learning stage to the data pre-processing stage, significantly enhancing training efficiency. The proposed TF-HNN can advance the field of hypergraph machine learning research by not only improving the efficiency of using observed hypergraph structures to solve downstream tasks, but also serving as a simple starting point for the development of future hypergraph machine learning models.
Limitation and future work. Learning tasks on hypergraphs can generally be defined at the node, hyperedge, and hypergraph levels. In this paper, we empirically test our TF-HNN on node-level and hyperedge-level tasks, with the main limitation being the lack of empirical results on hypergraph-level tasks. To our knowledge, designing task-specific modules to generate hypergraph-level features based on existing node features itself remains a challenging topic in the literature. Therefore, we consider the application of our TF-HNN to hypergraph-level tasks as future work. Another future direction of our work is to design a training-free module for explicit hyperedge-node importance learning.
@@ Appendix A Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix B Background of Hypergraph Message Passing
In this section, we provide proofs for the clique-based, star-based, tensor-based and line-based methods respectively.
@@
We first introduce some concepts about neighbours on hypergraphs and graphs: 1) The $k$-th hop neighbours of a node $v_{i}$ on a hypergraph are exactly $k$ hyperedges away from $v_{i}$; and 2) The $k$-hop neighbours of a node $v_{i}$ on a graph are all nodes with a distance of $k$ or less, while the $k$-th hop neighbours are exactly $k$ edges away from $v_{i}$. Before proving Proposition 3.3, we introduce some Lemmas.
@@
where $\delta(\cdot)$ is a function that returns 1 if $e_{k}$ connect $v_{i}$ and $v_{j}$ and returns 0 otherwise. We summarise the relationship between this edge weight and the probability of nodes $v_{i}$ and $v_{j}$ being in the same category as the following Lemma:
@@
By induction, for $L\in\mathbb{Z}_{+}$, $H^{v^{L}_{i}}_{0}=H^{v^{L}_{i}}_{2}$.
Now, we conduct the proof for the line-based models.
Let $H^{v^{L}_{i}}_{L}$ denote the entropy of the information contained by the $L$-hop neighbours of the line-nodes corresponding to $v_{i}$. According to previous papers in the GNN literature (Wu et al., 2020), the entropy of information used by an $L$-layer GNN to generate the features of line-nodes corresponding to $v_{i}$ is $H^{v^{L}_{i}}_{L}$. Based on the discussion in Appendix B and Appendix 3.1, we have $H^{v^{L}_{i}}_{0}=H^{v^{L}_{i}}_{L}$. By definition, the line-nodes corresponding to $v_{i}$ are only connected to each other and the line-nodes corresponding to the $1$-hop neighbours of $v_{i}$. Therefore, we have $H^{v^{L}_{i}}_{L}=H^{v^{L}_{i}}_{2}$. Accordingly, we have $H^{v^{L}_{i}}_{0}=H^{v^{L}_{i}}_{2}$. ∎
@@ Appendix I Details of Benchmarking Datasets
Table 9: The hyperlink prediction AUC (%) for HNNs and TF-HNN. The best result on each dataset is highlighted in bold font.
@@
Table 10: The training time for HNNs and TF-HNN in hyperlink prediction. The best result on each dataset is highlighted in bold font.
@@
Table 11: The node classification accuracy (%) for TF-HNN, HyperGCL, and HDSode. The best result on each dataset is highlighted in bold font.
@@
Table 12: The training time (s) for TF-HNN, HyperGCL, and HDSode. The best result on each dataset is highlighted in bold font.
@@
We adapt the algorithm from Chen & Liu (2023) for this purpose. For each (positive) hyperedge $e\in E$, we generate a corresponding negative hyperedge $f$, where $\alpha\times 100\%$ of the nodes in $f$ are from $e$ and the remaining are from $\𝑒V\textbackslash\{e\}$. Denote the negative hyperlink set as $F$. The number $\alpha$ controls the genuineness of the negative hyperlinks, i.e., higher values of $\alpha$ indicate that the negative hyperlinks are closer to the true ones. Additionally, we define $\beta$ to be the number of times of negative sampling which controls the ratio between positive and negative hyperlinks. In practice, we set $\alpha=0.5$ and $\beta=5$.
@@
Our code and data are available at: https://anonymous.4open.science/r/TF-HNN. To ensure the reproducibility of our results, we provide detailed hyperparameter settings used in our experiments. The configurations for the TF-HNN model on the hypergraph node classification task and the hyperlink prediction task are listed below.
Table 13: Hyperparameter settings for TF-HNN on the hypergraph node classification task.
@@
Table 14: Hyperparameter settings for TF-HNN on the hyperlink prediction task.

### v6

@@ 7 Conclusion
In this paper, we propose a novel model called TF-HNN. The key innovation of TF-HNN is an original training-free message passing module (TF-MP-Module) tailored for data on hypergraphs. We present both theoretical and empirical evidence demonstrating that TF-HNN can efficiently and effectively address hypergraph-related downstream tasks. To our knowledge, TF-HNN is the first model to shift the integration of hypergraph structural information from the model training stage to the data pre-processing stage, significantly enhancing training efficiency. The proposed TF-HNN can advance the field of hypergraph machine learning research by not only improving the efficiency of using observed hypergraph structures to solve downstream tasks, but also serving as a simple starting point for the development of future hypergraph machine learning models.
Limitation and future work. Learning tasks on hypergraphs can be defined at the node, hyperedge, and hypergraph levels. In this paper, we test our TF-HNN on node-level and hyperedge-level tasks, with the main limitation being the lack of results on hypergraph-level tasks. Designing task-specific modules to generate hypergraph-level features based on existing node features remains a challenging topic in the literature. Therefore, we consider the application of our TF-HNN to hypergraph-level tasks as future work. Finally, developing an automated and efficient hyperparameter search method for hypergraph neural networks presents a promising avenue for further enhancing model efficiency.
Acknowledgement
[…]
@@ Appendix A Impact Statements
This paper presents work that aims to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted.
@@ Appendix B Background of Hypergraph Message Passing
In this section, we provide proofs for the clique-based, star-based, tensor-based and line-based methods, respectively.
@@
We first introduce some concepts about neighbours on hypergraphs and graphs: 1) The $k$-th hop neighbours of a node $v_{i}$ on a hypergraph are exactly $k$ hyperedges away from $v_{i}$; and 2) The $k$-hop neighbours of a node $v_{i}$ on a graph are all nodes with a distance of $k$ or less, while the $k$-th hop neighbours are exactly $k$ edges away from $v_{i}$. We first introduce some Lemmas.
@@
where $\delta(\cdot)$ is a function that returns 1 if $e_{k}$ connects $v_{i}$ and $v_{j}$ and returns 0 otherwise. In this section, we discuss the relationship between this edge weight and the probability of nodes $v_{i}$ and $v_{j}$ being in the same category.
We summarise the relationship between this edge weight and the probability of nodes $v_{i}$ and $v_{j}$ being in the same category as the following Lemma:
@@
By induction, for $L\in\mathbb{Z}_{+}$, $H^{v^{L}_{i}}_{0}=H^{v^{L}_{i}}_{2}$. Now, we conduct the proof for the line-based models.
Let $H^{v^{L}_{i}}_{L}$ denote the entropy of the information contained by the $L$-hop neighbours of the line-nodes corresponding to $v_{i}$. According to previous papers in the GNN literature (Wu et al., 2020), the entropy of information used by an $L$-layer GNN to generate the features of line-nodes corresponding to $v_{i}$ is $H^{v^{L}_{i}}_{L}$. Based on the discussion in Appendix B and Proposition 3.1, we have $H^{v^{L}_{i}}_{0}=H^{v^{L}_{i}}_{L}$. By definition, the line-nodes corresponding to $v_{i}$ are only connected to each other and the line-nodes corresponding to the $1$-hop neighbours of $v_{i}$. Therefore, we have $H^{v^{L}_{i}}_{L}=H^{v^{L}_{i}}_{2}$. Accordingly, we have $H^{v^{L}_{i}}_{0}=H^{v^{L}_{i}}_{2}$. ∎
@@ Appendix I Details of Benchmarking Datasets
Table 9: The hyperlink prediction AUC (%). The best result is in bold font.
@@
Table 10: The training time (s) in hyperlink prediction. The best result is in bold font.
@@
Table 11: The node classification accuracy (%). The best result is in bold font.
@@
Table 12: The training time (s) in node classification. The best result is in bold font.
@@
We adapt the algorithm from (Chen & Liu, 2023) for this purpose. For each (positive) hyperedge $e\in E$, we generate a corresponding negative hyperedge $f$, where $\alpha\times 100\%$ of the nodes in $f$ are from $e$ and the remaining are from $Ve$. Denote the negative hyperlink set as $F$. The number $\alpha$ controls the genuineness of the negative hyperlinks, i.e., higher values of $\alpha$ indicate that the negative hyperlinks are closer to the true ones. Additionally, we define $\beta$ to be the number of times of negative sampling which controls the ratio between positive and negative hyperlinks. In practice, we set $\alpha=0.5$ and $\beta=5$.
@@
To ensure the reproducibility of our results, we provide detailed hyperparameter settings used in our experiments. The configurations for the TF-HNN model on the hypergraph node classification task and the hyperlink prediction task are listed below.
Table 13: Hyperparameter settings for TF-HNN in Table 3 and Table 5.
@@
Trivago 64 3 64 0.001 0.0 0.2 0.0
Table 14: Hyperparameter settings for TF-HNN in Figure 3 and Table 5.
@@
Trivago 1 3 64 0.0001 0.5 0.0
In practice, we conduct the hyperparameter search based on grid search. For the hypergraph node classification task listed in Table 3, we apply grid search to the following hyperparameters:
$\bullet$ The number of layers of TF-MP-Module: $\{1,2,4,8,16\}$.
$\bullet$ The $\alpha$ of TF-MP-Module: $\{0.05,0.1,0.15,0.2,0.25,0.3,0.35,0.4,0.45,0.5,0.55,0.6,0.65,0.7\}$.
$\bullet$ The number of layers of the node classifier: $\{3,5,7\}$.
$\bullet$ The hidden dimension of the node classifier: $\{128,256,512,1024\}$.
$\bullet$ The learning rate for node classifier: $\{1\times 10^{-4},2\times 10^{-4},3\times 10^{-4},4\times 10^{-4},5\times 10^{-4},6\times 10^{-4},7\times 10^{-4},8\times 10^{-4},9\times 10^{-4},1\times 10^{-3},5\times 10^{-3}\}$.
$\bullet$ The dropout rate for node classifier: $\{0.5,0.6,0.7,0.8,0.9\}$.
For node classification listed in Table 5, we apply grid search to the following hyperparameters:
$\bullet$ The number of layers of TF-MP-Module: $\{1,4,16,64\}$.
$\bullet$ The $\alpha$ of TF-MP-Module: $\{0.0,0.1,0.3,0.5\}$.
$\bullet$ The number of layers of the node classifier: $\{3\}$.
$\bullet$ The hidden dimension of the node classifier: $\{64\}$.
$\bullet$ The learning rate for node classifier: $\{1\times 10^{-3},1\times 10^{-4}\}$.
$\bullet$ The dropout rate for node classifier: $\{0.0,0.2,0.4\}$.
For hyperlink prediction listed in Figure 3, we apply grid search to the following hyperparameters:
$\bullet$ The number of layers of TF-MP-Module: $\{1,2,4,8,16,32,64\}$.
$\bullet$ The $\alpha$ of TF-MP-Module: $\{0.0,0.01,0.05,0.1,0.5\}$.
$\bullet$ The number of layers of the hyperedge predictor: $\{3\}$.
$\bullet$ The hidden dimension of the hyperedge predictor: $\{128,256\}$.
$\bullet$ The learning rate for node classifier: $\{1\times 10^{-4},2\times 10^{-4},3\times 10^{-4},4\times 10^{-4},5\times 10^{-4}\}$.
$\bullet$ The dropout rate for node classifier: $\{0.0,0.2,0.4,0.6,0.8\}$.
For hyperlink prediction listed in Table 5, we apply grid search to the following hyperparameters:
$\bullet$ The number of layers of TF-MP-Module: $\{1,2,4\}$.
$\bullet$ The $\alpha$ of TF-MP-Module: $\{0.0,0.1,0.3,0.5\}$.
$\bullet$ The number of layers of the hyperedge predictor: $\{3\}$.
$\bullet$ The hidden dimension of the hyperedge predictor: $\{64\}$.
$\bullet$ The learning rate for node classifier: $\{1\times 10^{-3},1\times 10^{-4}\}$.
$\bullet$ The dropout rate for node classifier: $\{0.0,0.2\}$.
The experiments on Trivago are highly time-consuming and need large GPU memory. We used a relatively small search space for hyperparameters in ours and the baselines, which are provided below.
Node classification on Trivago of TF-HNN:
$\bullet$ The number of layers of TF-MP-Module: $\{1,4,16,64\}$.
$\bullet$ The $\alpha$ of TF-MP-Module: $\{0.0,0.1,0.3,0.5\}$.
$\bullet$ The number of layers of the node classifier: $\{3\}$.
$\bullet$ The hidden dimension of the node classifier: $\{64\}$.
$\bullet$ The learning rate for node classifier: $\{1\times 10^{-4},1\times 10^{-3}\}$.
$\bullet$ The dropout rate for node classifier: $\{0.0,0.2,0.4\}$.
Node classification on Trivago of ED-HNN:
$\bullet$ The number of layers of ED-HNN: $\{1,2,4\}$ (larger values show the out-of-memory (OOM) issue on our server).
$\bullet$ The number of layers of $\hat{\phi}$, $\hat{\rho}$, and $\hat{\varphi}$, MLPs within message passing: $\{0,1,2\}$ (larger values show the out-of-memory (OOM) issue on our server).
$\bullet$ The number of layers of the classifier: $\{3\}$.
$\bullet$ The hidden dimension of $\hat{\phi}$, $\hat{\rho}$, and $\hat{\varphi}$: $\{64,128\}$ (larger values show the out-of-memory (OOM) issue on our server).
$\bullet$ The hidden dimension of the classifier: $\{64\}$.
$\bullet$ The learning rate for the model: $\{1\times 10^{-4},1\times 10^{-3}\}$.
$\bullet$ The dropout rate for for the model: $\{0.0,0.2,0.4\}$.
Node classification on Trivago of Deep-HGNN:
$\bullet$ The number of message passing layers: $\{1,4,16,64\}$.
$\bullet$ The $\alpha$: $\{0,0.1,0.3,0.5\}$.
$\bullet$ The $\lambda$: $\{0.1\}$.
$\bullet$ The message passing hidden dimension: $\{64,128,256\}$ (larger values show the out-of-memory (OOM) issue on our server).
$\bullet$ The number of layers of the classifier: $\{3\}$.
$\bullet$ The hidden dimension of the classifier: $\{64\}$.
$\bullet$ The learning rate for the model: $\{1\times 10^{-4},1\times 10^{-3}\}$.
$\bullet$ The dropout rate for for the model: $\{0.0,0.2,0.4\}$.
Hyperlink prediction on Trivago of ours:
$\bullet$ The number of layers of TF-MP-Module: $\{1,2,4\}$.
$\bullet$ The $\alpha$ of TF-MP-Module: $\{0.0,0.1,0.3,0.5\}$.
$\bullet$ The number of layers of the predictor: $\{3\}$.
$\bullet$ The hidden dimension of the predictor: $\{64\}$.
$\bullet$ The learning rate for the predictor: $\{1\times 10^{-4},1\times 10^{-3}\}$.
$\bullet$ The dropout rate for the predictor: $\{0.0,0.2\}$.
Hyperlink prediction on Trivago of ED-HNN:
$\bullet$ The number of layers of ED-HNN: $\{1,2,4\}$.
$\bullet$ The number of layers of $\hat{\phi}$, $\hat{\rho}$, and $\hat{\varphi}$, MLPs within message passing: $\{0,1,2\}$.
$\bullet$ The number of layers of the predictor: $\{3\}$.
$\bullet$ The hidden dimension of $\hat{\phi}$, $\hat{\rho}$, and $\hat{\varphi}$: $\{64,128\}$.
$\bullet$ The hidden dimension of the predictor: $\{64\}$.
$\bullet$ The learning rate for the model: $\{1\times 10^{-4},1\times 10^{-3}\}$.
$\bullet$ The dropout rate for for the model: $\{0.0,0.2\}$.
Hyperlink prediction on Trivago of Deep-HGNN:
$\bullet$ The number of message passing layers: $\{1,2,4\}$.
$\bullet$ The $\alpha$: $\{0,0.1,0.3,0.5\}$.
$\bullet$ The $\lambda$: $\{0.1\}$.
$\bullet$ The message passing hidden dimension: $\{64,128,256\}$.
$\bullet$ The number of layers of the classifier: $\{3\}$.
$\bullet$ The hidden dimension of the classifier: $\{64\}$.
$\bullet$ The learning rate for the model: $\{1\times 10^{-4},1\times 10^{-3}\}$.
$\bullet$ The dropout rate for for the model: $\{0.0,0.2\}$.